### v1

# Deep Reinforcement Learning-Based Active Flow Control of an Elliptical Cylinder: Transitioning from an Elliptical Cylinder to a Circular Cylinder and a Flat Plate

Abstract
This study investigates the effectiveness of Active Flow Control (AFC) technology supported by Deep Reinforcement Learning (DRL) in suppressing vortex shedding, reducing drag, and minimizing lift coefficients around elliptical cylinders. The aspect ratio ($Ar$) of the elliptical cylinder transitions from an ellipsoid ($Ar=2.0$) to a circular shape ($Ar=1.0$), and ultimately to a flat plate ($Ar=0$). We utilize the Proximal Policy Optimization (PPO) algorithm to precisely control the mass flow rates of synthetic jets located on both the upper and lower surfaces of the cylinder. The focus is on optimizing fluid dynamics and examining the scalability of these control techniques across different geometric configurations. Our research findings indicate that, for elliptical cylinders with aspect ratios between 1.75 and 0.75, the reduction in drag coefficient ranges from 0.9% to 15.7%, and the reduction in lift coefficient ranges from 95.2% to 99.7%. Notably, the DRL-based control strategy not only significantly reduces lift and drag, but also completely suppresses vortex shedding while using less than 1% of external excitation energy, demonstrating its efficiency and energy-saving capabilities. Additionally, for aspect ratios from 0.5 to 0, the reduction in drag coefficient ranges from 26.9% to 43.6%, and the reduction in lift coefficient from 50.2% to 68.0%. This reflects the control strategy’s significant reduction in both drag and lift coefficients, while also alleviating vortex shedding. This underscores the adaptability and potential of DRL-based AFC in managing complex fluid dynamics across diverse elliptical cylinder configurations.
preprint: AIP/123-QED
I Introduction
The development of active flow control technology has progressed from theoretical exploration to practical application, gradually becoming an important branch within the domains of fluid mechanics and engineeringCollis and Joslin (2004); Jahanmiri (2010); Awada, Younes, and Ilinca (2021). This technology aims to improve flow characteristics through active intervention in the flow field, such as reducing drag, controlling vortex shedding, enhancing mixing efficiency, or preventing flow separation, thereby enhancing system performance and efficiencyCollis and Joslin (2004); Wang et al. (2018). Initially, research in active flow control primarily focused on theoretical analysis and simple experiments, exploring how to influence fluid flow with external inputs like jets, vibrations, or magnetic fieldsJahanmiri (2010); Wang et al. (2018); Awada, Younes, and Ilinca (2021). With the advancement of Computational Fluid Dynamics (CFD) and increased computational power, researchers have been able to gain a deeper understanding of complex flow responses through numerical simulations and develop more effective control strategiesAram et al. (2018); Jain, Yeo, and Chopra (2010); Li et al. (2022). In recent years, breakthroughs in machine learning and deep learning technologies have led to revolutionary progress in the field of active flow controlBrunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Ren, Hu, and Tang (2020). Researchers have begun using intelligent algorithms, such as Deep Reinforcement Learning (DRL), to automatically optimize control strategies, achieving precise control over highly nonlinear and variable flow fieldsBrunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Garnier et al. (2021). These methods have not only improved the accuracy and real-time performance of control effects but also significantly expanded the application scope of active flow control technology, including but not limited to aerospace, automotive, energy, and environmental engineering sectorsVignon, Rabault, and Vinuesa (2023); Rabault et al. (2020); Garnier et al. (2021); Razdan and Shah (2022).
DRL is an avant-garde methodology that amalgamates the principles of deep learning with reinforcement learning to optimize decision-making through interaction with an environmentArulkumaran et al. (2017); François-Lavet et al. (2018); Wang et al. (2022a). Deep learning excels in autonomously extracting intricate features from voluminous datasets, while reinforcement learning concentrates on how to make optimal decisions based on these features, guiding the learning process through rewardsJaniesch, Zschech, and Heinrich (2021); LeCun, Bengio, and Hinton (2015). DRL leverages deep neural networks to represent the policies or value functions within reinforcement learning, thus empowering machines to tackle high-dimensional state spaces and complex decision-making challenges that were previously intractableArulkumaran et al. (2017); François-Lavet et al. (2018); Wang et al. (2022a). Notable for its robust non-linear modeling capabilities and adaptable learning mechanisms, DRL is capable of directly learning complex strategies from raw sensory data without the need for manually engineered featuresDargazany (2021). Moreover, DRL is inherently self-improving, continuously refining its strategies through ongoing interactions with the environmentWang et al. (2022a). This capacity for self-learning and adaptation bestows a significant advantage upon DRL, especially in resolving decision-making problems within dynamic settingsFrançois-Lavet et al. (2018); Wang et al. (2022a). The applications of DRL are extensive, encompassing domains such as gaming, autonomous vehicles, robotic control, natural language processing, and computer vision. In the gaming industry, breakthroughs like AlphaGo’s victory over world champions underscore DRL’s potentialGranter, Beck, and Papke (2017); in autonomous driving, DRL is employed for optimizing route planning and decision-makingSallab et al. (2017); in robotics, it enables robots to learn complex physical tasksZhao, Queralta, and Westerlund (2020). Additionally, DRL is being applied in scientific research areas like fluid mechanics and industrial optimization, showcasing its vast potential in managing complex systemsBrunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015).
In recent years, the application of DRL within the realm of fluid dynamics has seen a marked increase, with a primary focus on active flow control and shape optimizationBrunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Garnier et al. (2021). Active flow control technology is a key challenge in the fields of fluid mechanics and engineering, including goals such as drag reduction, prevention of flow separation, control of vortex shedding, and enhancement of mixing efficiencyWang et al. (2018). Traditional control methodologies often rely on a priori knowledge and empirical rules, whereas DRL offers a novel pathway for autonomously learning optimal control strategies from dataCollis and Joslin (2004); Jahanmiri (2010). In the context of drag reduction applications, DRL is utilized to autonomously adjust the operation of control devices (e.g., jets, vibratory mechanisms) to minimize the fluidic drag around objects such as cylindrical bodies or car modelsGarnier et al. (2021); Razdan and Shah (2022). For vortex shedding control, DRL is capable of learning how to dynamically adjust control parameters to stabilize the flow field and reduce vibrations and noise caused by vortex sheddingVignon, Rabault, and Vinuesa (2023); Rabault et al. (2020). In the domain of shape optimization, DRL facilitates the automatic discovery and application of effective shape adjustments to achieve optimization goals by intelligently exploring the design space and learning optimal design strategiesBhola et al. (2023). With the continued expansion of computational resources and advancements in algorithms, the application of DRL in fluid dynamics is set to broaden further.
Fluid mechanics problems are characterized by a high degree of nonlinearity and multiscalar nature, involving high-dimensional state spaces and complex system dynamicsPozorski and Wacławczyk (2020); Reichstein et al. (2019). These challenges are often intractable for traditional mathematical models, demanding sophisticated control strategies, particularly when addressing phenomena such as unstable flows, turbulence, and multiphase flows, which require consideration of a wide range of spatiotemporal scalesSubramaniam (2020). The suitability of DRL methods for these kinds of fluid mechanics issues primarily rests on several key features: Firstly, DRL can handle the high-dimensional and complex state spaces intrinsic to fluid dynamics, effectively capturing the intricate dynamics presentXie et al. (2023). Through the function approximation capabilities of deep neural networks, DRL is adept at processing fluid mechanics systems with numerous variables and parameters, discerning patterns and regularities hidden within dataZhao et al. (2024). Secondly, DRL’s interactive and adaptive learning approach allows it to self-tune and optimize strategies in response to dynamic and uncertain fluid environments. Traditional control methods often struggle with such dynamicsBhansali et al. (2023); however, DRL models, through their interaction with the environment and trial-and-error processes, can independently learn and refine strategies to meet diverse fluidic conditions and task requirementsHusen, Chaudary, and Ahmad (2022). Moreover, the long-term reward mechanism intrinsic to DRL allows for the consideration of future impacts when formulating strategies. Control and optimization in fluid mechanics often necessitate foresight into long-term effectsVinuesa et al. (2022). DRL’s focus on long-term rewards enables models to balance short-term gains with future outcomes, thereby better addressing control and optimization challenges within fluid mechanicsViquerat et al. (2022); Vinuesa (2024). In conclusion, the attributes of DRL render it highly suitable for addressing the control and optimization challenges in fluid mechanics.
Rabault et al.Rabault et al. (2019) pioneered the application of advanced DRL algorithms in active flow control. The author uses the DRL agent to control the mass flow rate of the synthetic jets on the upper and lower sides of the cylinder, effectively reducing the resistance around the cylinder by 8% and mitigating the detachment of the rear vortex. Rabault et al.; Tang et al.; Heess et al.; Ren, Rabault, and Tang; Jia and XuRabault et al. (2019); Tang et al. (2020); Heess et al. (2017); Ren, Rabault, and Tang (2021); Jia and Xu (2024a) applied the Proximal Policy Optimization (PPO) algorithm to active flow control (AFC) in the flow around a cylinder. The training results consistently showed a reduction in drag of 8% around the cylinder. Li and ZhangLi and Zhang (2022) investigated the flow stability and sensitivity in confined cylinder wakes under various blockage ratios. They used the analysis of flow stability and structure sensitivity to guide the manipulation of the flow field by DRL agents. Wang et al.Wang et al. (2022b) utilized an AFC method based on the PPO algorithm to reduce the drag on a cylindrical body by approximately 8%, and applied the Soft Actor-Critic (SAC) algorithm to decrease the drag around a square cylinder by about 13.6%. Jia and XuJia and Xu (2024b) employed the SAC algorithm within a DRL framework to achieve AFC around a square cylinder, effectively modulating the flow dynamics across varying Reynolds numbers. Wang et al.Wang et al. (2022c) used DRL methods for active control of flow through elliptical cylinders. Their training results prove that DRL can successfully learn active control strategies under four angles of attack and various long and short axis ratios. Fan et al.Fan et al. (2020) was the first to apply DRL to active flow control experiments, optimizing power efficiency by adjusting the rotational speed of two small cylinders behind the main cylinder. Ren, Wang, and TangRen, Wang, and Tang (2021) proposed a novel active flow control strategy to hide the hydrodynamic characteristics of bluff bodies, including strong shear forces and periodic shedding vortices, to avoid predators.
The application of AFC strategies based on DRL has been widely validated in various bluff body flow scenarios for their effectiveness in controlling vortex shedding and reducing drag. Among these, studies on flow around constrained cylinders under different blockage ratios represent the most thoroughly investigated areaRabault et al. (2019); Tang et al. (2020); Li and Zhang (2022); Wang et al. (2022b); Jia and Xu (2024a). By contrast, research related to flow around elliptic cylinders has seen relatively less exploration within this domain. Due to their unique shape, elliptic cylinders influence the development of separation points and boundary layers differently during flow, potentially leading to Hopf bifurcation and periodic vortex shedding at lower Reynolds numbers. These flow characteristics offer opportunities in aerospace, marine engineering, energy engineering, and modern architectural design for improving lift-to-drag ratios, reducing vortex-induced vibration, enhancing energy efficiency, and increasing structural stability. Therefore, an in-depth investigation of the flow characteristics around elliptic cylinders not only enhances our understanding of fluid dynamics but also promotes performance optimization and innovation in engineering design.
In summary,despite significant advancements in controlling flow around circular cylinders through DRL algorithms, the study of flow around elliptical cylinders presents a vast research landscape due to its more complex flow characteristics and control challenges. AFC based on DRL algorithms introduces a novel methodology that leverages learning from flow field information to intelligently adjust control strategies, aiming at reducing drag and suppressing vortex shedding oscillations. This research is dedicated to examining the impact of various $Ar$ of elliptical cylinders on flow stability and sensitivity to structural parameters. Building upon this, it explores the application of DRL-based AFC in managing flow around elliptical cylinders by utilizing the derived physical insights. By establishing a robust reinforcement learning environment and defining clear reward mechanisms, this study aims to identify and implement effective flow control strategies. The aim of this study is to enhance the flow stability around an elliptical cylinder using AFC technology based on DRL algorithm, serving as a reference for flow control around similarly complex-shaped objects.
The structure of this paper is organized as follows: Section II introduces the problem description and provides a comprehensive overview of Deep DRL and its framework for application in AFC. Section Section III presents the main results and analysis. Finally, Section IV concludes the paper by summarizing the key findings and discussing their implications.
II Problem description and methodology
II.1 Numerical simulation
Model configuration
The present study investigates the flow around two-dimensional elliptical cylinders with their major axis oriented perpendicular to the flow direction, as depicted in Fig. 1(a). The major axis of the constrained elliptical cylinder (denoted 2a) remains fixed, while the shape of the elliptical cylinder is changed by changing the value of 2b. The Reynolds number is calculated based on the length of the major axis, formulated as $Re=U(2a)/\nu$, where $U$ represents the velocity of the incoming uniform flow and $\nu$ denotes the kinematic viscosity. The aspect ratio ($Ar$) is defined as the ratio of the semi-minor axis ($b$) to the semi-major axis ($a$), expressed as $Ar=b/a$. For ease of comparison with circular cylinders, we define $2a$ as $D$, representing the diameter of an equivalent circular cylinder. Consequently, an $Ar$ of 0 designates a flat plate normal to the flow, while an $Ar$ of 1 denotes a circular cylinder. This study investigates a series of variations ranging from an elliptical cylinder $(Ar=2.0)$ to a circular cylinder $(Ar=1.0)$, and further to a flat plate $(Ar=0)$, as shown in Fig. 1(c).
Figure 1: (a) Definition of the problem, including relevant dimensions. (b) Establishment of boundary conditions. (c) This study examines an elliptical cylinder with varying aspect ratios: a normal flat plate corresponding to a cylinder with an aspect ratio of 0, a special case of a cylinder with an aspect ratio of 1, and an elliptical cylinder with an aspect ratio of 2.
Computational domain and coordinate system
An elliptical cylinder is placed within a computational domain bounded by a rectangular perimeter, where the diameter of the elliptical cylinder is defined as D, and the major axis length (2a) of the elliptical column is equivalent to D. The computational domain extends from -2D in front of the cylinder or elliptical column to 20D behind it in the flow direction, which is the direction of fluid flow. In the cross-flow direction, perpendicular to the flow direction, the domain ranges from -2D to 2.1D. Within this computational domain, a Cartesian coordinate system is established with the center of the cylinder or elliptical column as the origin O, where the direction of flow is defined along the x-axis, and the axis perpendicular to the flow direction is the y-axis. The configuration of the synthetic jets is depicted in Fig. 1(b), specifically positioned at two directions on the cylindrical or elliptical column, $\theta_{1}=90^{\circ}$ and $\theta_{2}=270^{\circ}$, representing symmetrical locations of the jet on either side of the object’s surface. The width of the jet opening is fixed at $\omega=10^{\circ}$.
Governing equations
Let $\Omega\subset\mathbb{R}^{nd}$ and $(0,T)$ be the spatial and temporal domains, respectively, where $nd$ is the number of space dimensions, and let $\Gamma$ denote the boundary of $\Omega$. The spatial and temporal coordinates are denoted by $x$ and $t$. This work investigates the dynamics of incompressible viscous fluid flow. The Navier–Stokes equations governing incompressible fluid flow are
$\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot(\nabla\mathbf{u})=-\nabla p+Re^{-1}\Delta\mathbf{u}\quad\text{on}\quad\Omega\times(0,T),$ (1a)
$\nabla\cdot\mathbf{u}=0\quad\text{on}\quad\Omega\times(0,T).$ (1b)
Here, $\bm{u}=(U,V)$ represents the velocity vector, where $U$ and $V$ are the velocity components along the $x$- and $y$-axes, respectively. $p$ is the thermodynamic pressure. The $Re$ is calculated based on the length of the major axis, formulated as $Re=U(2a)/\nu$, where $U$ represents the velocity of the incoming uniform flow and $\nu$ denotes the kinematic viscosity.
Boundary conditions
The upstream boundary of the computational domain, designated as $\Gamma_{\text{in}}$, is characterized by a parabolic velocity profile oriented in the $x$-axis, where the boundary conditions are specified such that the streamwise velocity component $U$ is equal to the mean velocity $U$ with a value of 1, and the transverse velocity component $V$ is set to zero. On the downstream boundary, denoted as $\Gamma_{\text{out}}$, a Neumann-type boundary condition is prescribed for the velocity field, ensuring that the stress vector is nullified. On the upper and lower boundaries $\Gamma_{\text{w}}$ , a slip-wall boundary condition is employed, which stipulates that both the velocity component perpendicular to these boundaries and the tangential stress vector component along them are assigned a value of zero. The synthetic jets boundaries are denoted by $\Gamma_{i}(i=1,2)$. The velocity within the jets adheres to a parabolic distribution, with the sign of the jet velocity indicative of fluid suction or injection. The design of the synthetic jets ensures mass flow balance, as expressed by the condition $V_{\Gamma_{1}}=-V_{\Gamma_{2}}$, to maintain mass conservation in the system. Except for the jet, the non-slip solid wall boundary condition $\Gamma_{\text{w}}$ is adopted for the rest of the cylinder.
In compliance with the above requirements, the system is considered with the following boundary conditions
$-\rho\mathbf{n}\cdot\mathbf{p}+Re^{-1}(\mathbf{n}\cdot\nabla\mathbf{u})$ $=0\quad\text{on }\Gamma_{out},$ (2)
$\mathbf{u}$ $=0\quad\text{on }\Gamma_{w},$
$\mathbf{u}$ $=U\quad\text{on }\Gamma_{in},$
$\mathbf{u}$ $=f_{Q_{i}}\quad\text{on }\Gamma_{i},\quad i=1,2.$
Here, $U$ is the inflow velocity profile. $f_{Q_{i}}$ represents the radial velocity profiles simulating the suction or injection of fluid by the jets. The specific formulas of $U$ and $f_{Q_{i}}$ will be given below.
At the inlet $\Gamma_{\text{in}}$, the inflow velocity $U$ along the $x$-axis is prescribed by a parabolic velocity profile in the form,
$U=U_{m}\frac{(H-2y)(H+2y)}{H^{2}}.$ (3)
Here, $U_{m}$ is the maximum velocity magnitude of the parabolic profile, and $H=4.1D$ represents the total height of the rectangular domain.
The functions $f_{Q_{i}}$ represent radial velocity profiles that simulate the suction or injection mechanisms of the jets. These profiles are selected to ensure a smooth transition to the no-slip boundary conditions on the surfaces $\Gamma_{W}$ of the cylinder. Specifically, $f_{Q_{i}}$ is defined as $A(\theta;Q_{i})(x,y)$, with the modulation relying on the angular coordinate $\theta$ as illustrated in Fig. 2(b). For the jet characterized by a width of $\omega$ and centered at $\theta_{0}$, which is positioned on the cylinder with radius $R$, the modulation is established as follows:
$A(\theta;Q)=Q\frac{\pi}{2\omega R^{2}}\cos\left(\frac{\pi}{\omega}(\theta-\theta_{0})\right).$ (4)
Solver details
In this study, the incompressible flow solver is based on the open-source CFD package OpenFOAM, as described by Jasak et al.Jasak (2009); Jasak et al. (2007) OpenFOAM is a widely used and validated CFD software package that provides robust numerical algorithms for solving the Navier-Stokes equations. The solver in OpenFOAM utilizes the finite volume method to discretize the computational domain into a mesh made up of control volumes. To ensure numerical stability, the time step is chosen as $\Delta t$ = 0.0005.
Grid system
Taking the computational domain of an elliptical cylinder with an $Ar$ of 1 as an exemplar, the domain undergoes discretization utilizing a synergistic approach that incorporates both structured (quadrilateral) and unstructured (triangular) meshes. As delineated in Fig. 2, this process results in the domain being segmented into 18,484 mesh elements. Herein, the vicinity of the elliptical cylinder is discretized with multiple layers of quadrilateral meshes, whereas the remaining regions are discretized using triangular meshes. Fig. 2(a) reveals the global discretization scheme of the computational domain, highlighting the comprehensive approach to mesh division. Fig. 2(b) offers insight into the mesh partitioning adjacent to the elliptical cylinder, with the multi-layered quadrilateral mesh around the cylinder being significantly finer compared to the triangular mesh. Fig. 2(c) illustrates the multi-layered quadrilateral mesh division around the elliptical cylinder and details the integration strategy between structured and unstructured meshes. Lastly, Fig. 2(d) provides a closer inspection of the mesh division surrounding the elliptical cylinder, emphasizing the precision of the discretization near the structure. The mesh division method for ellipses with $Ar$ other than 1 follows the same approach as that for $Ar=1$.
Figure 2: Numerical simulation of water flowing through an elliptical cylinder, close-up view of the mesh at $Ar=1$. The computational domain is discretized into 18,484 mesh elements, incorporating both triangular and quadrilateral meshes. (a) Discretization of the computational domain is viewed from a global perspective. (b) A comprehensive display of the mesh around the elliptical cylinder. (c) Presentation of the multi-layered quadrilateral meshes surrounding the elliptical cylinder. (d) A magnified view of the quadrilateral meshes in the vicinity of the elliptical cylinder.
Quantities of interest
The lift coefficient ($C_{L}$) and drag coefficient ($C_{D}$) are defined as
$C_{L}=\frac{F_{L}}{0.5\rho\overline{U}^{2}D},\quad C_{D}=\frac{F_{D}}{0.5\rho\overline{U}^{2}D}.$ (5)
Here, $F_{L}$ and $F_{D}$ represent the lift and drag forces integrated on the surface of the elliptical cylinder, respectively, and $\rho$ is the fluid density. The Strouhal number $(St)$ is used to describe the characteristic frequency of oscillatory flow phenomena and is defined as follows:
$St=\frac{f\cdot D}{U}.$ (6)
Where $f_{s}$ is the shedding frequency calculated based on the periodic evolution of the $C_{L}$. $D$ is the characteristic length of the elliptical cylinder, and $\overline{U}$ is the mean velocity of the upstream flow.
Grid independence
Taking the case with an $Ar$ of 1 as an example, this study explores mesh convergence and validates the numerical method utilized. Tab. 1 presents the simulation results using three meshes with differing resolutions, enabling a comparative analysis of the maximum drag coefficient $C_{D,\max}$, the mean of the maximum drag coefficient $C_{D,\text{mean}}$, the maximum lift coefficient $C_{L,\max}$, and the $St$. Furthermore, the results of the numerical simulations are compared with the findings of Rabault et al. Rabault et al. (2019) and Schäfer et al. Schäfer et al. (1996) for validation. The simulations conducted using a coarse mesh revealed significant deviations in the drag and lift coefficients when compared to the results by Rabault et al. Rabault et al. (2019), questioning the accuracy of such numerical simulations. On the other hand, the fine mesh, characterized by a high density of grid points, considerably slows down the training speed for reinforcement learning, presenting a challenge for efficient simulation execution. Our aim is to perform numerical simulations with the least number of mesh elements without compromising on computational accuracy. Consequently, the discretization strategy of the main mesh was adopted for subsequent numerical simulations and reinforcement learning training. This strategy not only satisfies the computational accuracy requirements but also keeps the mesh element count to a manageable level. Comparison of the computational results from the main mesh with those obtained by Rabault et al. Rabault et al. (2019). shows very little difference, indicating a high level of accuracy. Furthermore, the accuracy of the main mesh’s results lies within the upper and lower bounds provided by standard benchmarksSchäfer et al. (1996), confirming that the computational precision of the main mesh is adequate for further training purposes. The above analysis ensures the best balance between numerical accuracy and computational efficiency. The methodology for examining mesh dependency for ellipses with $Ar$ other than 1 is consistent with the procedure established for $Ar=1$.
Table 1: Grid Independence Test for an Elliptical Cylinder with $Ar=1$ at $Re=100$.
Configuration Mesh resolution Mesh $C_{D,max}$ $C_{D,mean}$ $C_{L,max}$ $St$
Rabault et al.Rabault et al. (2019) - 9262 3.205 -
Schäfer et al.Schäfer et al. (1996) - - 3.220–3.240 - 0.990–1.010 0.295–0.305
Coarse 10540 3.242 3.224 1.052 0.304
$Ar=1$ Medium 18484 3.225 3.205 0.990 0.300
Fine 25624 3.228 3.207 0.992 0.301
II.2 Deep Reinforcement learning
Deep learning
Deep learning, a subset of machine learning, employs multi-layer artificial neural networks to model and learn intricate nonlinear relationships. Fig. 3(a) demonstrates how deep learning captures the representation of input data across multiple layers of a neural network, including the input layer, several hidden layers, and the output layer. Each layer is composed of numerous neural units, with information transmitted between nodes via connection weights. The fundamental concept behind deep learning is its ability to autonomously learn data representations through successive layers. Utilizing the backpropagation algorithm, the network autonomously discerns the transformation process from raw data to sophisticated, high-level abstract features, eliminating the need for manual feature extraction.
(a) DL
(b) RL
(c) DRL
Figure 3: The basic architecture of deep learning, reinforcement learning, and deep reinforcement learning.
Reinforcement learning
In the domain of Reinforcement Learning, the interaction between an agent and its environment is the core mechanism by which learning is facilitated. As illustrated in Figure 3(b), this interaction proceeds in a cyclical pattern. At each timestep $t$, the agent observes the current state $s_{t}$ of the environment, which contains all the relevant information necessary for making decisions. Based on this observation, the agent selects and executes an action $a_{t}$, according to its policy $\pi$. This policy, which is essentially a strategy that maps states to actions, may function deterministically or probabilistically. After the action is executed, the environment transitions to a new state $s_{t+1}$ and awards the agent with an immediate reward $r_{t+1}$. This reward, quantified as a numerical value, evaluates the action’s efficacy in progressing towards the agent’s objectives. Subsequently, the agent assimilates this feedback—comprised of both the reward and the new state—to refine its policy, employing RL algorithms designed to optimize the accumulation of long-term rewards.
Deep Reinforcement learning
Deep learning, a specialized subset of machine learning, excels at identifying patterns in extensive datasets through its multi-layer neural network architecture. This approach enables computers to derive complex representations directly from raw data, leading to revolutionary advances in fields like image recognition, natural language processing, and predictive analytics. DRL marries the potent perceptual capabilities of deep learning with the decision-making and optimization capabilities of reinforcement learning, effectively overcoming their individual limitations. Deep learning addresses the challenge of processing original, high-dimensional input data, which is problematic for traditional reinforcement learning, while reinforcement learning provides a structured decision-making framework and a goal-oriented learning mechanism that enhances the capabilities of deep learning. As depicted in Figure 3(c), a deep learning network transforms the intricate state of the environment into a compressed and efficient feature representation. This transformation yields a more manageable and interpretable state space for reinforcement learning algorithms. In DRL, reinforcement learning leverages these high-level features to learn and optimize decision-making policies. These policies are typically parameterized by deep neural networks, implying that the selection of actions is influenced not only by the current observation of the environmental state but also by the learned weights within the network.
II.3 PPO algorithm
PPO is an advanced DRL algorithm suitable for solving reinforcement learning problems in discrete and continuous action spaces. The core idea is to improve performance by maximizing approximate policy optimization while maintaining the stability of policy updates. In PPO, the agent interacts with the environment, collects empirical data, and uses this data to construct an objective function. The objective function takes into account dominance estimates from empirical data and is used to measure the direction of policy improvement. In the PPO algorithm, the agent interacts with the environment to generate a series of trajectories ($\tau$). A trajectory is a sequence of state-action pairs produced by the agent following the current policy interacting with the environment until reaching a terminal state or a predetermined maximum number of steps. Symbolically, a trajectory $\tau$ can be represented as:
$\tau=(s_{0},a_{0},r_{1},s_{1},a_{1},r_{2},s_{2},\ldots,s_{T-1},a_{T-1},r_{T},s_{T}),$ (7)
where, $s_{t}$ represents the state at time step $t$; $a_{t}$ represents the action at time step $t$; $r_{t+1}$ represents the immediate reward obtained after taking action $a_{t}$; $T$ represents the length of the trajectory; $s_{0}$ is the initial state, and $s_{T}$ is the terminal state. The cumulative reward $R(\tau)$ of a trajectory $\tau$ is obtained by summing the rewards $r_{t}$ obtained at each time step $t$:
$R(\tau)=\sum_{t=0}^{T}r_{t},$ (8)
The objective function $J(\pi_{\theta})$ in the PPO algorithm aims to maximize the expected cumulative reward obtained by following the policy $\pi_{\theta}$. Mathematically, this objective function is expressed as:
$J(\pi_{\theta})=\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)],$ (9)
where $J(\pi_{\theta})$ represents the expected return under policy $\pi_{\theta}$, with the expectation taken over trajectories sampled from the policy distribution $\pi_{\theta}$. The term $R(\tau)$ denotes the total reward accumulated along a trajectory $\tau$. The policy gradient $\nabla_{\theta}J(\pi_{\theta})$, indicating the direction of steepest ascent in the objective function $J(\pi_{\theta})$, is computed using the policy gradient theorem. The gradient of the expected return is expressed as the expectation of the gradient of the log-probability of trajectories weighted by the return:
$\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}_{\tau\sim\pi_{\theta}}[\nabla_{\theta}\log P(\tau\mid\theta)R(\tau)],$ (10)
where $\nabla_{\theta}\log P(\tau\mid\theta)$ represents the gradient of the log-probability of trajectories with respect to the policy parameters $\theta$. This computation involves sampling trajectories from the policy distribution $\pi_{\theta}$ and weighting the log-probability gradients by their respective returns. The policy parameters $\theta$ are updated iteratively using gradient ascent, with the update rule:
$\theta_{k+1}=\theta_{k}+\alpha\nabla_{\theta}J(\pi_{\theta}),$ (11)
where $\theta_{k}$ represents the policy parameters at the $k$-th iteration, $\alpha$ is the learning rate controlling the step size, and $\nabla_{\theta}J(\pi_{\theta})$ denotes the policy gradient.
In the PPO algorithm, the trajectories generated by the agent are instrumental in shaping the policy update process. The objective function, $J(\pi_{\theta})$, is designed to enhance the agent’s performance by maximizing the expected cumulative reward under the policy $\pi_{\theta}$. This optimization involves iteratively updating the policy parameters $\theta$ using the policy gradient $\nabla_{\theta}J(\pi_{\theta})$, while maintaining stability and preventing substantial policy shifts. The policy gradient is estimated using a collection of trajectories, $\mathcal{D}=\{\tau_{i}\}_{i=1,\ldots,N}$, each generated by the agent interacting with the environment according to the policy. The policy gradient can be calculated as follows:
$\hat{g}=\frac{1}{|\mathcal{D}|}\sum_{\tau\in\mathcal{D}}\sum_{t=0}^{T}\nabla_{\theta}\log\left(\pi_{\theta}(a_{t}\mid s_{t})\right)R(\tau),$ (12)
where $|\mathcal{D}|$ represents the number of trajectories in $\mathcal{D}$, and $T$ is the length of each trajectory. This formulation offers a simplified approach to computing the policy gradient. Moreover, PPO commonly employs a clipped surrogate objective function to facilitate more stable policy updates. This objective is expressed as:
$L^{CLIP}(\theta)=\frac{1}{|\mathcal{D}|}\sum_{\tau\in\mathcal{D}}\sum_{t=0}^{T}\min\left(r_{t}(\theta)A_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)A_{t}\right),$ (13)
where $r_{t}(\theta)$ is the likelihood ratio between the old and new policies, $A_{t}$ denotes the advantage function, and $\epsilon$ is a hyperparameter that defines the clipping range. The inclusion of this clipped surrogate objective aids in moderating the extent of policy updates, thereby stabilizing the training process.
In summary, the PPO algorithm presents a robust framework for the training of deep reinforcement learning agents. It operates through an iterative process of updating policy parameters aimed at maximizing the expected cumulative reward. Crucially, PPO maintains stability and mitigates the risk of substantial policy deviations. Employing policy gradients along with clipped surrogate objectives, PPO ensures effective and stable training across a diverse array of reinforcement learning tasks. This methodology not only enhances the efficiency of learning but also contributes significantly to the consistency and reliability of the training process.
II.4 DRL-Enhanced Active Flow Control
This section outlines the design of a framework utilizing DRL algorithms to address active flow control problems. Firstly, it is necessary to define two fundamental elements in DRL: the agent and the environment. The agent can be designed as algorithms suitable for solving continuous action control problems, such as PPO agent or SAC agent. The CFD simulation of the flow around an elliptical cylinder is chosen as the environment for the agent to interact with. Rabault et al.Rabault et al. presented the main structure of the custom environment class, encouraging readers to delve into the detailed content. This framework design applies deep reinforcement learning methods to solve the active flow control problem, enabling the agent to continuously learn and optimize its behavior through interaction with the environment to achieve the control goal. Fig. 4 illustrates the fundamental framework of DRL, the computational process of a single episode during training, and the formulas utilized.
Figure 4: DRL leverages the concurrent execution of multiple environments within a parallel computational framework. Each iteration consists of two distinct phases: exploration and optimization. During the exploration phase, the agent engages with the environment via its current policy $\pi$, executing actions $a_{t}$, receiving states $s_{t}$ and rewards $r_{t}$, and noting terminal states. Upon the conclusion of the iteration, the parallel computation across $n$ environments yields a set of trajectories $\{\tau_{i}\}_{i=1}^{n}$, with each trajectory $\tau_{i}$ comprising the sequence of states, actions, and rewards as $\tau_{i}=\{(s_{0},a_{0}),(s_{1},a_{1},r_{1}),\ldots,(s_{T},a_{T},r_{T})\}$, where $(s_{t},a_{t},r_{t})$ represents the state, action , and reward at time-step $t$, respectively, for the $i$-th environment. In the optimization phase, the agent samples a minibatch of trajectories from the experience replay buffer to compute the policy gradient $\nabla_{\theta}J(\pi_{\theta})$. The parameters $\theta$ of the policy network are updated using gradient ascent, i.e., $\theta\leftarrow\theta+\alpha\nabla_{\theta}J(\pi_{\theta})$, where $\alpha$ denotes the learning rate. This process iterates, with the agent deploying the updated policy $\pi_{\theta^{\prime}}$ to interact with the environment, computing the advantage function $A(s_{t},a_{t})$, and recalculating the policy gradient for subsequent updates of $\theta$. Through iterative refinement, the policy network incrementally enhances its action selection strategy within the environment. The formula delineates the update rule applied to the policy network during the optimization phase.
  • State $s_{t}$: Traditional state spaces are often too expansive for effective policy learning, requiring operations in partially observable environments. As a result, an agent’s performance heavily relies on the quality and relevance of the data it can observe. The impact of partial observability of states $s_{t}$ on control results was mentioned in Viquerat et al.Viquerat et al. (2022)’s research.Integrating instantaneous velocity or pressure information from CFD simulations to characterize the environmental state $s_{t}$, the DRL agent can effectively perceive and analyze the current flow conditions. To achieve the control objectives of reducing drag and lift around the elliptical cylinder and controlling the shedding of vortices in the wake field, multiple probes are strategically placed around the elliptical cylinder and within its wake field. This method is advantageous for gathering comprehensive data pertaining to the flow dynamics around the cylinder and subsequent wake characteristics. The design of probes placement in this study was based on the design principles articulated by Rabault et al.Rabault et al. (2019), which advocate strategic positioning of probes around the cylinder and within the wake region of the cylinder.In their studies, Jia and XuJia and Xu (2024a) and Wang et al.Wang et al. (2022b) all referred to the probes arrangement scheme proposed by Rabault et al.Rabault et al. (2019).
  • Action $a_{t}$ : The control action involves the blowing and suction of two synthetic jets on the elliptical cylinder. The flow rates of the two synthetic jets are constrained such that $Q_{\text{lower}}+Q_{\text{upper}}=0$. Therefore, the RL agent manipulates the velocity of the upper synthetic jets, while the lower synthetic jets has the same velocity but in the opposite direction.The velocity of the synthetic jets is set to be less than 2% of the inlet mean velocity. This restriction is imposed to prevent the agent from taking shortcuts and using excessively high jet velocities to achieve drag reduction, which contradicts the fundamental principle of active flow control: energy efficiency. Furthermore, each control action lasts for 50 time steps, termed as a viscous action. Allowing faster actions could harm overall performance as they need time to influence the environment. Ensuring smooth transitions between current and updated actions is crucial. Let $a_{t}$ and $a_{t+1}$ represent the action magnitudes at consecutive time steps. To ensure a gradual transition between these actions and mitigate abrupt changes in jet velocity, a smoothing function $S$ can be defined mathematically. The function $S$ interpolates between the previous action magnitude $V_{\Gamma_{1},T_{i-1}}$ and the target action magnitude $a$ at time step $i$, resulting in an updated value $V_{\Gamma_{1},T_{i}}$. This smoothing function can be expressed as:
    $S(V_{\Gamma_{1},T_{i}},a,V_{\Gamma_{1},T_{i-1}})=V_{\Gamma_{1},T_{i}}+\beta\cdot(a-V_{\Gamma_{1},T_{i-1}}),$ (14)
    where $V_{\Gamma_{1},T_{i}}$ represents the current value at time step $i$, $\beta$ is a coefficient determining the extent of adjustment towards the target action $a$, and $V_{\Gamma_{1},T_{i-1}}$ is the value at the previous time step $i-1$. This function ensures a smooth transition between consecutive actions, with the parameter $\beta$ controlling the smoothness of the transition.
  • Reward $r_{t}$ : This study incorporates the average drag coefficient (without control), the time-averaged value of the drag coefficient, and the time-averaged value of the lift coefficient (averaged over the duration of the action). The primary objective is to minimize the drag while maintaining a low lift coefficient; hence, the weight of the drag coefficient in the reward function is set to 1, while the lift coefficient is assigned a weight $w$, typically chosen to be 0.1 or 0.2. This hyperparameter $w$ is determined through tuning. Including a constant (the average drag coefficient without control) in the reward function helps guide it towards positive rewards, which positively affects the convergence speed of the learning curve. The reward function $f(\cdot)$ translates the drag and lift coefficients into a numerical reward assessing the agent’s behavior. The specific reward function takes the following form:
    $r_{T_{i}}=C_{D,0}-\left(C_{D}\right)_{T_{i}}-\omega\left|\left(C_{L}\right)_{T_{i}}\right|,$ (15)
    $C_{D,0}$ represents the baseline drag coefficient, serving as a reference point for drag reduction. $(C_{D})_{T_{i}}$ denotes the drag coefficient at time step $T_{i}$, with the objective of minimizing this value relative to the baseline. $(C_{L})_{T_{i}}$ signifies the lift coefficient at time step $T_{i}$, whose absolute value is penalized to mitigate lift forces that may destabilize the flow around the bluff body. $\omega$ is a weighting factor that quantifies the trade-off between minimizing drag and controlling lift fluctuations.
  • Parameter. Time granularity in the DRL framework is crucially set based on the fluid dynamics problems studied and significantly affects agent performance and learning task complexity. Precise adjustments of this hyperparameter, informed by prior research and our ongoing strategy refinement, are vital for achieving optimized control outcomes. Each action is applied for a duration of 0.025 seconds (Corresponds to 25 numerical simulation time steps), and each episode consists of 100 control actions. The maximum total time for an episode is defined as $T_{\max}=T_{100}=2.5$ nondimensional time units. The shedding frequency of vortices around the elliptical cylinder in this study ranges from 0.3 to 0.4, corresponding to 6.25 to 8.33 vortex shedding cycles per episode. The length of episodes is intentionally chosen to encompass multiple shedding cycles, facilitating effective learning of the control algorithm by the agent through observation and adaptation of flow dynamics over several cycles.
II.5 High-Performance Computing in DRL and CFD
The DRL training and numerical simulation in this study were conducted on a high-performance computing(HPC) system equipped with an Intel® Xeon® Platinum 8358 CPU, operating at 2.60GHz and comprising a total of 64 cores distributed evenly across two sockets, each containing 32 cores. Numerical simulations were performed using the eighth version of the open-source software platform, OpenFOAM® version 8.Jasak et al. (2007) The deployment of deep reinforcement learning algorithms was based on the Tensorforce library.Schaarschmidt, Kuhnle, and Fricke (2017) The ANN definitions and gradient descent algorithms were facilitated using the TensorFlow open-source library.Abadi et al. (2016) For DRL implementation, agent was loaded using the Tensorforce platform with the PPO algorithm, and the environment was constructed using the Gym interface.Brockman et al. (2016) Fig. 4 illustrates the use of CFD calculations as the environment for DRL training and shows the effects of using 60 parallel CFD environments to accelerate the training process. This parallel strategy was inspired by the research of Rabault and KuhnleRabault and Kuhnle (2019a) and Wang et al.Wang et al. (2022b), and also incorporated insights from the work of Wang et al.Wang et al. (2022b) For a more comprehensive understanding, readers are encouraged to consult their works and the references therein.
III Results
III.1 Baseline Flow
Vortex shedding phenomenon
Figure 5: Instantaneous Vorticity Contours for Flow Around an Elliptic Cylinder at $Re=100$, Varying from $Ar=2$ to $Ar=0$. Purple and orange indicate positive and negative vorticity, respectively. In all the depicted vorticity contours, purple represents the clockwise vortices shed from the upper side of the elliptical cylinder, while orange represents the counterclockwise vortices shed from the lower side.
Fig. 5 presents the vortex shedding patterns in the wake region formed around elliptical cylinders with varying $Ar$. The subfigures (a) to (j) depict the vortex structures that have reached a fully developed state, illustrating the progression from an aspect ratio of 0 to 2. At a Reynolds number of 100, around elliptical cylinders with $Ar$ varying from 0 to 2, persistent unstable vortices are observed at the computational domain’s outlet, indicating the continuation of fluid instability along the outflow direction. However, as the fluid progresses downstream, there is a diminishing trend in the peak vortex intensity. This attenuation of vorticity is predominantly attributed to viscous diffusion effects prevalent in low Reynolds number regimes. This is consistent with the study by Johnson, Thompson, and HouriganJohnson, Thompson, and Hourigan (2004). Between $Ar$ of 2 and 1.25, in addition to the recirculation bubbles downstream of the cylinder, two additional recirculation bubbles form upstream near the wall. As the $Ar$ decreases further, the confined flow exhibits only one recirculation bubble. Similar to observations by Sahin and OwensSahin and Owens (2004), under the influence of blockage ratio and Reynolds number, two additional recirculation bubbles appear on the solid sidewalls of the computational domain. At specific Reynolds numbers, this will lead to another Hopf bifurcation at $Re_{\text{c2}}$.
For an elliptical cylinder with an $Ar=2$, the wake field does not exhibit vortex shedding, and the zero vorticity contour lines show slight fluctuations. This suggests that despite the overall stability of the wake, there is still minor flow instability. As the $Ar$ decreases to 1.75 and 1.5, the wake field begins to exhibit more pronounced vortex shedding phenomena. The vortices become more compact, and there is increased interaction between positive and negative vortices through shearing, stretching, and detachment processes. As the $Ar$ decreases to 1, transforming the cylinder into a perfect circle, the vortex structure becomes more regular, forming a typical von Kármán vortex street. This "standard aligned BvK street" is characterized by a series of regularly spaced vortices that alternately form on both sides of the cylinder. As the $Ar$ decreases from $Ar=0.75$ to $Ar=0.25$, the recirculation area gradually diminishes, leading to a more compact vortex structure and an increase in shedding frequency. At $Ar=0.1$, for extremely low $Ar$, the recirculation area further decreases, resulting in an increase in both the frequency and intensity of vortex shedding. At $Ar=0$, when the cylinder degenerates into a flat plate, particularly in the zero-vortex region behind the ellipse, the acceleration of vortex shedding becomes more pronounced. This indicates an increase in flow instability and irregularity in vortex size and shape.
Fluctuation velocity
Strykowski and SreenivasanStrykowski and Sreenivasan (1990) demonstrated that the introduction of minor disturbances in a targeted region can alter vortex formation, effectively reducing or controlling vortex shedding. Similarly, our study employs small perturbations to disturb the flow field with the aim of decreasing drag and suppressing vortex shedding. Furthermore, the judicious incorporation of fluid physics knowledge within the control agents can streamline their architecture without compromising performance, as exemplified by Pastoor et al.Pastoor et al. (2008), where phase control in drag reduction of a bluff body is primarily founded upon the understanding of decoupled shear layer development and wake processes. In this study, the strategic placement of probes was thoughtfully engineered to align with the dynamic and controllable aspects of the flow field. Particular attention was paid to its inherent instability and pulsatility, ensuring that the probes could effectively capture the critical fluctuations necessary for a comprehensive analysis of the flow dynamics. The research conducted by Protas and WesfreidProtas and Wesfreid (2002) explores the dynamics of the total drag coefficient, $C_{D}$, which is defined by the equation $C_{D}=C_{D}^{b}+C_{D}^{0}$. Here, $C_{D}^{b}$ represents the base drag coefficient, a static value consistent across specific $Re$, encapsulating the inherent drag associated with the baseline flow. Conversely, the correctable component $C_{D}^{0}$ reflects the influence of oscillatory flow phenomena on the drag coefficient, providing an avenue for adjustment through strategic manipulation of Bénard–von Kármán vortices. In our study, the instantaneous velocity field is segmented into the mean velocity field and the fluctuating velocity field. We hypothesize that probes equipped to detect these pulsatile and oscillatory dynamics can leverage targeted minor disturbances to effectively neutralize these instabilities. Such interventions are anticipated to facilitate the suppression of vortex shedding, thereby enhancing the control of flow behaviors.
Figure 6: For the case of $Ar=1$, the distribution of probes is depicted across various velocity fields: (a) Instantaneous velocity, (b) Fluctuating velocity.
We delineates the instantaneous and fluctuating velocity fields behind an elliptical cylinder during full vortex development, as illustrated in Fig. 6. The strategic placement of probes is predicated on the objectives of reducing drag and lift around the elliptical cylinder while simultaneously suppressing vortex shedding. Probes are positioned around and downstream of the elliptical cylinder, specifically within the vortex shedding region. The probes around the elliptical cylinder aim to diminish the drag and lift coefficients integral to our control objectives. Downstream, probes are strategically deployed in the wake to facilitate the extraction of flow field information from the regions of maximum fluctuating velocity, where intense turbulent motion typically occurs due to vortex instability and enhanced turbulent energy. The probe placement in the cylinder’s wake encompasses the entire recirculation area, which is critical for capturing the most unstable and intensely turbulent information in the wake region. This configuration plays a significant role in reducing the fluctuating components of the drag coefficient, suppressing vortex shedding, and stabilizing the wake flow. In Fig. 6, an ellipse with an $Ar$ of 1 serves as the prototype. The methodology for designing probe placements for other $Ar$ adheres to the same principles established for $Ar=1$.
III.2 Effectiveness of AFC
This section describes the performance of AFC using the DRL algorithm, which manipulates the magnitude of synthetic jets to achieve targeted flow control. Within a flow configuration at $Re=100$, AFC is applied to elliptical cylinders with $Ar$ ranging from 2 to 0, as well as to cylindrical and flat plate configurations. The evaluation of the control’s effectiveness and performance is based on several criteria: comparisons of the velocity fields before and after control, reductions in drag coefficient, variations in lift coefficient, the actions (magnitude of the synthetic jets velocities), and the learning curves observed during the training process.
Figure 7: Snapshots of the velocity magnitude illustrating the effect of the active flow control. Left side of each row: baseline simulation. Left side of each row: result with active flow control. (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$; (f) $Ar=0.75$.
Fig. 7 and Fig. 8 provide flow field snapshots of elliptical cylinders under various $A$, encompassing both baseline (uncontrolled) and controlled scenarios. This section primarily describes the flow structures observed in the baseline cases. For elliptical cylinders with $Ar=2$, the baseline flow field inherently displays high stability, with no vortex shedding observed. Consequently, the AFC did not produce a noticeable effect. As the aspect ratio decreases to $Ar=1.75$, vortices begin to detach alternately from both sides of the elliptical cylinder at reduced frequencies and intensities. These detaching vortices organize into a relatively orderly and stable vortex street. By applying AFC based on DRL, the vortex shedding in the wake of the elliptical cylinder is completely suppressed, resulting in a stabilized wake field. At an aspect ratio of $Ar=1.5$, an alternating vortex structure forms behind the elliptical cylinder. These structures possess certain strengths and frequencies, yet the extent of vortex shedding remains comparatively mild. Further reduction in the aspect ratio to $Ar=1.25$ leads to a more pronounced and frequent shedding of vortical structures from the wake, indicating a transition towards more significant flow instability. The recirculation bubble downstream of the elliptical cylinder diminishes, oscillates, and becomes unstable. When the aspect ratio is $Ar=1.5$ and $Ar=1.25$, a stable, regular alternating vortex structure forms behind the elliptical cylinder. After the application of AFC, the shedding of vortices behind the elliptical cylinder is completely suppressed, transitioning the wake field into a stable flow structure. At $Ar=1$ and $Ar=0.75$, the area of the recirculation bubbles formed behind the elliptical cylinder decreases, leading to more prominent and frequent vortex shedding within the wake, indicating a transition towards significant flow instability. Following the implementation of AFC, the alternating vortex street behind the elliptical cylinder is transformed into a stable wake field.
Fig. 7 depicts the instantaneous velocity fields both baseline and post-implementation of control measures across a range of $Ar$ from 2 to 0.75 for elliptical cylinders. At $Ar=2$, the baseline wake field is notably stable. However, when the $Ar$ is reduced below 1.75, alterations in structural parameters provoke enhanced vortex shedding behind the ellipse, characterized by escalating frequency and intensity as the $Ar$ diminishes. The efficacy of AFC is underscored through the deployment of a control strategy, which effectively attenuates vortex shedding behind the elliptical cylinder for $Ar=1.75$, 1.5, 1.25, 1, and 0.75. This intervention demonstrates the significant capabilities of the DRL-based AFC technology in manipulating vortex shedding within the wake fields of elliptical cylinders across varying $Ar$, thereby enhancing the flow stability of the configurations under study.
Figure 8: Snapshots of the velocity magnitude illustrating the effect of the active flow control. Left side of each row: baseline flow. Left side of each row: result with active flow control. (a) $Ar=0.5$; (b)$Ar=0.25$; (c) $Ar=0.1$; (d) $Ar=0$.
As the $Ar$ is further reduced to 0.5, 0.25, and 0.1, eventually transitioning to a flat plate configuration, the vortex shedding phenomena behind the elliptical cylinder become more frequent and intense. It is evident that decreasing the $Ar$ leads to a significant increase in the circulation entering the wake, as the flow velocity increases when water moves around the top of the cylinder. This observation aligns with the findings of Thompson et al.Thompson et al. (2014). This underscores that as the $Ar$ of an elliptical cylinder gradually diminishes towards zero, it induces an enhancement in the flow instability of the wake field, consequently leading to the release of more energy in the form of vortices. This dynamic highlights the critical role of structural parameters in influencing the energy distribution and stability of the flow field as it interacts with varying geometrical configurations.
However, when utilizing the AFC technique to manipulate the flow field, the control outcomes differ from those observed at $Ar=0.75$ and higher $Ar$; vortex shedding behind the elliptical cylinder is not completely suppressed. At $Ar=0.5$, analysis of the instantaneous velocity field reveals that the recirculation area behind the elliptical cylinder is significantly elongated, and the shedding of vortices is notably subdued. Nevertheless, the flow field has not fully stabilized, as there remains a continual detachment of alternating vortex structures from behind the elliptical cylinder, resembling elongated vortex formations. At $Ar=0.25$, the recirculation bubble behind the elliptical cylinder is similarly elongated as seen at $Ar=0.5$, but with a notable difference: the bubble now exhibits significant oscillations. In the controlled wake field, the phenomenon of flow separation following fluid passage around the elliptical cylinder and subsequent reattachment is also observable. Despite these dynamics, the instability of the downstream velocity field has been relatively stabilized compared to the baseline state, indicating the effectiveness in delaying vortex shedding. At $Ar=0.1$ or 0, compared to the baseline condition, the control of the flow field has primarily delayed rather than completely suppressed vortex shedding behind the elliptical cylinder. The recirculation bubble behind the cylinder is elongated, yet it exhibits significant oscillations. This indicates that while the AFC has extended the duration before vortex detachment, it has not fully stabilized the vortex structures within the wake.
In summary, at higher $Ar$ ($Ar=0.75$, 1, 1.25, 1.5, 1.75), the DRL-based control strategy can fully suppress vortex formation, thereby stabilizing the wake field. However, at lower $Ar$ ($Ar=0.5$, 0.25, 0.1, 0), the effectiveness of the control strategy is limited to mitigating vortex shedding within the wake field to some extent. It delays the vortex detachment process and enlarges the recirculation area behind the elliptical cylinder, but oscillations in the recirculation bubble persist, and the downstream flow field continues to exhibit signs of instability.
Figure 9: Comparison of $C_{D}$ and $C_{L}$ before and after control.(a) and (b) $Ar=2$; (c) and (d) $Ar=1.75$; (e) and (f) $Ar=1.5$; (g) and (h) $Ar=1.25$; (i) and (j) $Ar=1$; (k) and (l) $Ar=0.75$.
Fig. 9 presents a comparative analysis of the $C_{D}$ and $C_{L}$ before and after control intervention over $Ar$ ranging from 2 to 0.75. At an $Ar$ of 2, both $C_{D}$ and $C_{L}$ are stable, indicating that AFC does not have any effect. However, for $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, upon activating synthetic jets control, there is an immediate, sharp response in $C_{D}$, which quickly descends to a value near its minimum. After half the total control duration, $C_{D}$ stabilizes and remains constant at this minimal value. Similarly, $C_{L}$ experiences significant fluctuations at the start of control but stabilizes near zero after half the total control duration. The reductions in $C_{D}$ and $C_{L}$ at different $Ar$ are detailed in Tab. 2. These observations substantiate the effectiveness of the active flow control strategy underpinned by DRL algorithms. The activation of control swiftly alters the periodic behavior of both $C_{D}$ and $C_{L}$, facilitating the stabilization of $C_{D}$ and $C_{L}$ in the subsequent stages. This demonstrates the capability of the DRL-based active flow control approach to significantly reduce the drag around an elliptical cylinder to its lowest feasible value and diminish the lift to zero.
Fig. 10 illustrates the $C_{D}$ and $C_{L}$ for both baseline and controlled flow conditions around an elliptical cylinder at $Ar$ of 0.5, 0.25, 0.1, and 0. Under AFC, at $Ar=0.5$ and $Ar=0.25$, both $C_{D}$ and $C_{L}$ are reduced compared to the baseline flow, yet they do not achieve the stable state observed at $Ar=0.75$ and higher. Initially, both the $C_{D}$ and the $C_{L}$ experienced abrupt changes. However, they did not fully stabilize in the subsequent period. In the latter half of the control duration, $C_{D}$ and the $C_{L}$ exhibited fluctuations, but both the amplitude and frequency of these fluctuations were lower than those observed in the baseline flow. At $Ar=0.1$ and $Ar=0$, the controlled flow experiences a sharp decrease in the drag coefficient at the onset of control, followed by oscillations that persist at a lower frequency and amplitude than the baseline flow. Although complete stability is not achieved, the average value of the $C_{D}$ is significantly reduced. The $C_{L}$ undergoes an initial sharp decrease and approaches near zero, and after half the total duration of control, it oscillates at a lower frequency.
Figure 10: Snapshots of the velocity magnitude illustrating the effect of the active flow control. Left side of each row: baseline flow. Left side of each row: result with active flow control. (a) $Ar=0.5$; (b)$Ar=0.25$; (c) $Ar=0.1$; (d) $Ar=0$.
Figure 11: Action, which is the mass flow rate of the synthetic jets. (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$. (f) $Ar=0.75$; (g) $Ar=0.5$; (h) $Ar=0.25$; (i) $Ar=0.1$; (j) $Ar=0$.
Fig. 11 illustrates the time history curve of the action (the mass flow rate of synthetic jets), When performing AFC using DRL for flow around an elliptical cylinder with $Ar$ between 0 and 2. When $Ar=2$, the AFC is ineffective, and the value of the action, pertaining to the control input, remains close to zero. From $Ar=1.75$ to $Ar=0.75$, during the initial phases of control, the agent selects larger action values, corresponding to higher mass flow rates of synthetic jets. This indicates a deliberate strategy to assert flow control by introducing significant external energy inputs. As control progresses, the agent gradually diminishes the action values, thereby reducing the mass flow rate of the synthetic jets. Subsequently, in the mid to later stages, the agent’s action values converge towards zero, corresponding to a near-zero mass flow rate of synthetic jets. This behavior signifies the agent’s optimization towards an energy-saving state, maintaining flow characteristics with minimal control input. When $Ar=0.5$ and $Ar=0.25$, the value of the action initially surges but does not stabilize near zero in the later stages. For $Ar=0.1$ and $Ar=0$, the action value remains in a state of periodic fluctuation.
Overall, the results presented in Fig. 7 and Fig. 9 illustrate that the DRL-based control strategy exhibits energy efficiency and effectiveness within the range of elliptical $Ar=0.75$ to $Ar=1.75$. By progressively decreasing the mass flow rate of synthetic jets, the strategy achieves flow control, reduces drag and lift forces, and does so with reduced external energy input. However, at $Ar=0.5$ and $Ar=0.25$, the energy efficiency of the AFC strategy diminishes. At $Ar=0.1$ and $Ar=0$, the external energy utilized by the control strategy is relatively high compared to the inlet velocity, which will be detailed in Tab. 2.
Figure 12: The learning curve of PPO algorithm training when $Ar=2$ to 0. (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$. (f) $Ar=0.75$; (g) $Ar=0.5$; (h) $Ar=0.25$; (i) $Ar=0.1$; (j) $Ar=0$.
In Fig. 12, we observe the learning curve during training of flow around an elliptical cylinder with different $Ar$. At $Ar=2$, the reward function value never reaches a positive figure, indicating that the DRL training did not achieve its objectives. For $Ar=1.75$ and $Ar=1.5$, the reward function initially increases from negative to positive values and then after 4000 episodes converges near the maximum reward value. At $Ar=1.25$, despite some initial fluctuations, the training reaches a relatively stable state after 3000 episodes and eventually settles into a stable plateau. For $Ar=1$ and $Ar=0.75$, there are significant initial fluctuations with the learning curve rising swiftly, indicating rapid progress in policy learning. After 2000 episodes, the training demonstrates a stable convergence trend, with minor fluctuations in the reward value between 2000 and 5000 episodes. This shows that the PPO algorithm has successfully learned an effective strategy, exhibiting commendable convergence and stability in performance. At $Ar=0.5$, the speed of training convergence begins to slow, trending towards convergence after 4000 episodes, with the reward values stabilizing between 4000 and 10000 episodes. At $Ar=0.25$, training approaches convergence after 8000 episodes, reaching a total of 20000 episodes, with a mid-to-late plateau marking stable convergence. At $Ar=0.1$ and $Ar=0$, the training convergence is extremely slow, with signs of convergence only appearing after nearly 12000 episodes.
Overall, the DRL algorithm exhibits rapid initial convergence and reaches a stable plateau during the mid-to-late stages for $Ar=1.25$, 1 and 0.75, showcasing excellent convergence and stability throughout the training process. Training also achieves stable convergence within 5000 episodes for $Ar=1.5$ and $Ar=1.75$. However, for $Ar=0.5$, 0.25, 0.1 and 0, as the flow phenomena become increasingly complex, the control challenge intensifies, leading to a noticeable slowdown in convergence speed. This indicates that for flow scenarios with higher degrees of instability, the convergence performance of DRL training is compromised, requiring a larger number of training iterations to reach stable convergence.
Tab. 2 presents the effectiveness of active flow control techniques based on DRL algorithms, targeting elliptical cylinders with $Ar$ ranging from 0.1 to 2. It meticulously details the values of $C_{D}$ and $C_{L}$ before and after the application of control, along with their respective rates of change. Additionally, it describes the magnitude of the mass flow rate of the synthetic jets utilized for control and its ratio relative to the inflow rate. Initially, at $Ar=2$, the wake field stability behind the elliptical cylinder is observed to be relatively high, resulting in no significant alteration in the drag experienced by the elliptical cylinder. When $Ar$ is reduced to 1.75, there is a noticeable decrease in drag around the elliptical cylinder by 0.9%. With further reduction in $Ar$, the rate of drag reduction progressively increases. In their study, Protas and Wesfreid (2002) elucidated that the drag coefficient consists of two components: (i) the invariant drag ($C_{D}^{b}$), which remains constant at a given Reynolds number; and (ii) the amendable drag correction ($C_{D}^{0}$), associated with oscillatory flow phenomena within the fluid, representing the additional resistance arising from fluid oscillations or fluctuations. This portion of drag can be adjusted through external interventions, indicating the influence of unstable flow regions and the potential for control. In this research, the flow field is controlled by utilizing "external excitation," such as synthetic jets, to mitigate the additional resistance caused by instability, thereby optimizing the overall fluid dynamic performance. As the $Ar$ of the elliptical cylinder decreases, the pulsations, oscillations, and instabilities within the wake intensify, thereby enhancing the controllable component of the drag.
Table 2: The control effect of elliptical cylinder under DRL-based AFC.
$Ar$ Re $\overline{C}_{D,\text{Baseline}}$ $\overline{C}_{D,\text{Controlled}}$ Drag Reduction (%) $\overline{C}_{L,\text{Baseline}}$ $\overline{C}_{L,\text{Controlled}}$ Lift Reduction (%) $\overline{a}$ $\overline{a}$ ratio (%)
0 100 7.679 4.331 43.6 1.513 0.605 60.0 0.521 52.1
0.1 100 7.812 4.876 37.6 1.196 0.596 50.2 0.499 49.9
0.25 100 6.503 4.167 35.9 1.707 0.579 66.1 0.126 12.6
0.5 100 4.837 3.533 26.9 1.937 0.619 68.0 0.049 4.9
0.75 100 3.792 3.197 15.7 1.603 0.076 95.2 0.010 1.0
1 100 3.207 2.951 8.0 1.022 0.011 98.9 0.001 0.1
1.25 100 2.918 2.804 3.9 0.541 0.006 98.9 0.006 0.6
1.5 100 2.765 2.707 2.1 0.302 0.002 99.4 0.009 0.9
1.75 100 2.707 2.684 0.9 0.165 0.001 99.7 0.007 0.7
2 100 2.684 2.684 0.0 0.016 0.016 0 0 0
  • a. $\overline{C}_{D,\text{Baseline}}$ and $\overline{C}_{L,\text{Baseline}}$ represents the time-averaged drag coefficient of the elliptical cylinder once the baseline flow has fully developed.
  • b. $\overline{C}_{D,\text{Controlled}}$, $\overline{C}_{L,\text{Controlled}}$, and $\overline{a}$ denote the time-averaged values under stable conditions after AFC is applied, excluding the transient response phase.
  • c. $\overline{a}$ ratio (%) represents the ratio of the mass flow rate of the synthetic jets to the flow rate from the inlet.
In Tab. 2, for elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the $C_{D}$ reduction rates are 0.9%, 2.1%, 3.9%, 8.0%, and 15.7% respectively, while the $C_{L}$ reduction rates exceed 95% in each case. These results highlight the efficacy of the DRL-based AFC technology in effectively suppressing both $C_{D}$ and $C_{L}$ within the $Ar$ range of 1.75 to 0.75, demonstrating the efficiency of the DRL-based intelligent control strategy. However, as $Ar$ decreases to 0.5, 0.25, 0.1, and 0, even though the $C_{D}$ reduction rates significantly increase to 26.9%, 35.9%, 37.6%, and 43.6%, the $C_{L}$ reduction rates are 68.0%, 66.1%, 50.2%, and 60.0% respectively. This indicates that while the DRL-based control strategy achieves significant reductions in $C_{D}$, its effect on suppressing $C_{L}$ is limited and does not completely eliminate $C_{L}$ around the elliptical cylinder. As the $Ar$ further decreases, the AFC technique based on the DRL algorithm does not fully suppress vortex shedding behind the elliptical cylinder, but merely delays the shedding process. The reduction in $C_{L}$ coefficients aligns with the flow snapshots presented in Fig. 8, showing that vortex shedding still occurs, albeit with controlled frequency and intensity. These phenomena suggest that at very low $Ar$ values, the control strategy faces challenges in managing more complex and unstable flow conditions.
[…]
The control strategies derived from DRL-based AFC technology exhibit greater complexity compared to conventional harmonic forcing implemented in previous works (Bergmann, Cordier, and BrancherBergmann, Cordier, and Brancher (2005)), highlighting the value of employing ANN as controllers. Typically, approximately 93% of the drag induced by the shedding of vortices is suppressed by the discovered control laws, with minimal jet intensities required for drag reduction. For elliptical cylinders with aspect ratios $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the magnitude of the mass flow rate injected by the jets normalized by the mass flow rate of the mainstream intersecting with the cylinder is within 1%.
III.3 Analysis of the Controlled Flow
Fig. 13 illustrates the suppression of vortex shedding in the wake of elliptical cylinders under the influence of active flow control based on DRL. The elliptical cylinders examined have $Ar$ of 1.75, 1.5, 1.25, 1.0, and 0.75, scenarios in which vortex shedding can be completely suppressed. At the initial moment of control (T1), vortices in the wake of the elliptical cylinders are fully developed. During this phase, vortices are observed detaching from the cylinder at a consistent frequency and intensity, forming a Kármán vortex street in the downstream flow. As the control process advances (T2), external excitation methods, such as synthetic jets, are introduced to suppress the vortices in the wake of the elliptical cylinders. The objective of the control is to alter the characteristics of the flow field to weaken or eliminate the formation and persistence of vortices. With the progression of the control measures, a snapshot of the flow field is captured at the conclusion of the control (T3). At this juncture, vortices within the wake of the elliptical cylinders are observed to be suppressed, demonstrating the effectiveness of the DRL-based active flow control strategy in stabilizing complex fluid dynamics.
(a) $Ar=2$
(b) $Ar=1.75$
(c) $Ar=1.5$
(d) $Ar=1.25$
(e) $Ar=1$
(f) $Ar=0.75$
Figure 13: When $Ar=2$, 1.75, 1.5, 1.25, 1, 0.75, the vortex shedding of the elliptical cylinder can be completely suppressed. The streamwise snapshots of the wake flow field behind an elliptical cylinder, where the vortex shedding process is suppressed, feature key moments designated as T1, T2, and T3. These represent the instances of the onset of control (when the vortex is fully developed), during the control process, and at the conclusion of control, respectively.
In the study of elliptical cylinders, the phenomenon of vortex shedding and its control are crucial for understanding the structure of flow fields. For elliptical cylinders with different $Ar$, the behavior of the flow fields shows significant differences. At $Ar=2$, the instantaneous vorticity maps at times T1, T2, and T3 demonstrate a high degree of stability in the flow field, with no evident vortex shedding observed. For elliptical cylinders with $Ar$ values of 1.75, 1.5, and 1.25, at T1, alternating vortices form behind the cylinder, influenced by the elliptical shape effect, which results in lower intensity and frequency of vortex shedding. By T2, with the initiation of control, synthetic jets quickly intervene in the regular vortex shedding process, leading to a delay in vortex detachment and the formation of "elongated vortex structures." The original vortices are significantly elongated and narrowed, and the instantaneous recirculation bubbles are also elongated longitudinally and tend towards stability. By T3, the original vortex shedding has been completely suppressed, with no new vortices forming, and the instantaneous recirculation bubbles reach their maximum extent and remain stable, demonstrating the effectiveness of the control strategy.
At $Ar=1$ and $0.75$, during the T1 phase, a standard Bénard–von Kármán vortex street forms behind the elliptical cylinder, characterized by a series of regularly spaced vortices. These vortices alternately shed from both sides of the cylinder, creating a regular pattern of vortex arrangement. Moving into the second phase (T2), similar to cases with higher aspect ratios ($Ar=1.25$, 1.5, 1.75), the recirculation bubbles behind the cylinder are elongated longitudinally, and the originally shed vortices are also elongated and narrowed, with shearing interactions occurring between them. By the third time point (T3), the vortices behind the elliptical cylinder are completely suppressed, and the regions of positive and negative vorticity in the flow field nearly achieve symmetry, demonstrating high stability.
(a) $Ar=0.5$
(b) $Ar=0.25$
(c) $Ar=0.1$
(d) $Ar=0$
Figure 14: When $Ar=0.5$, 0.25,0.1 and 0, vortex shedding from the elliptical cylinder is not entirely suppressed. The streamwise snapshots of the wake flow field behind the elliptical cylinder capture key moments designated as T1, T2, and T3. These correspond to the initiation of control (when the vortex is fully developed), the early phase of the control process, the mid-to-late phase of the control process, and the termination phase of control, respectively.
Fig. 14 delineates scenarios where the suppression of vortex shedding from elliptical cylinders is not entirely effective for aspect ratios $Ar$ of 0.5, 0.25, 0.1, and 0. At T1, the vortex shedding behind the cylinder is fully developed, T2 represents the control phase, and T3 provides a snapshot of the flow field after control cessation. For $Ar=0.5$ and $Ar=0.25$, During the T1 phase, vortices are observed detaching from the cylinder with certain frequency and intensity, forming a Kármán vortex street in the downstream wake. In the T2 phase, as control commences, the recirculation bubble is noted to elongate longitudinally to some extent, with the recirculation area appearing relatively symmetric and stable. The intervention of synthetic jets delays the detachment of vortices, which are stretched and deformed due to shear with the side walls. By T3, the initially generated vortices are significantly elongated and deformed, exiting the computational domain, yet the recirculation bubble behind the cylinder remains unstable and continues to oscillate, leading to new vortex shedding. The emergence of new irregular vortex shedding highlights persistent instability behind the elliptical cylinder. For elliptical cylinders with $Ar=0.5$ and $Ar=0.25$, the DRL-based control strategy does not completely suppress vortex shedding. The instability of the wake field is mitigated, demonstrating partial efficacy of the control strategy in manipulating flow dynamics, although full stability is not achieved.
For $Ar=0.1$ and $Ar=0$, during T1, the vortex shedding frequency and intensity from the cylinder are more severe compared to $Ar=0.5$, with enhanced flow instability. Multi-level vortex shedding occurs, with primary and secondary vortices differing in shape, frequency, and intensity. At T2, vortices behind the cylinder are elongated and delayed, and the recirculation area significantly enlarges. The instability of the recirculation area manifests as asymmetry, swinging, and oscillation phenomena. By T3, although the original shedding vortices are elongated and expelled from the computational domain, new, unstable, and irregular shedding occurs. Shed vortices swinging to the side walls are sheared, stretched, and undergo deformation, resulting in irregularly shaped vortices. This highlights the persistent challenges in managing flow dynamics and instability at lower aspect ratios.
IV Conclusions
In this study, we investigate the use of DRL-based AFC to suppress vortex shedding, drag, and lift coefficients of an elliptical cylinder with a Reynolds number of 100. The $Ar$ of the elliptical cylinder ranges from an ellipse ($Ar=2.0$) to a circle ($Ar=1.0$), and finally to a flat plate ($Ar=0$). We employ the PPO algorithm to precisely control the mass flow rates of synthetic jets located on the upper and lower surfaces of the elliptical cylinder. The primary objective of this work is to gain further insights into the control capabilities of deep reinforcement learning algorithms for complex flow systems. The main findings of this study can be summarized as follows:
  • We evaluated the impact of $Ar$ on flow control around elliptical cylinders. When the $Ar$ is 2, the baseline flow is highly stable, and AFC shows negligible effect. However, for aspect ratios of 1.75, 1.5, 1.25, 1, and 0.75, the reduction rates in the drag coefficient around the cylinder are respectively 0.9%, 2.1%, 3.9%, 8.0%, and 15.7%, while the reduction rates in lift coefficient are 99.7%, 99.4%, 98.9%, 98.9%, and 95.2%. These data indicate that the AFC strategy based on DRL is capable of significantly minimizing the cyclically fluctuating drag coefficient to minimal values and stabilizing it, while also maintaining the cyclically oscillating lift coefficient near zero. This demonstrates the efficacy of DRL in AFC applications, effectively reducing the lift and drag around elliptical cylinder with $Ar$ ranging from 0.75 to 1.75. Velocity cloud diagrams reveal that the original periodically shedding vortices transition to a stable flow field with no vortex shedding after flow control. Instantaneous vorticity maps during the control process show that with the activation of synthetic jets, the vortices shedding downstream of the elliptical cylinder are progressively elongated and squeezed out of the computational domain. The originally unstable recirculation area behind the cylinder also transforms into a stable, symmetrical recirculation bubble, ceasing the production of new vortices. These results underscore the robust adaptability of DRL-based AFC across different elliptical cylinder shapes ($Ar$ between 0.75 and 1.75), capable of learning optimal control strategies directly through interaction with the environment to achieve a reduction in drag and lift coefficients around the cylinder. Moreover, for elliptical cylinders with aspect ratios of 1.75, 1.5, 1.25, 1, and 0.75, the mass flow rate ratios of synthetic jets relative to the inlet flow rate are 0.7%, 0.9%, 0.6%, 0.1%, and 1.0% respectively. This underscores that the flow control strategy efficiently utilizes less than 1% of external excitation energy, exemplifying its energy-saving capability and fully satisfying the low-energy consumption requirements essential for effective flow control.
  • When the $Ar$ of the elliptical cylinder decreases to 0.5, 0.25, 0.1, and 0, the reduction rates of the drag coefficient and lift coefficient are 26.9%, 35.9%, 37.6%, and 43.6%, respectively, while the corresponding reduction rates for the lift coefficient are 68.0%, 66.1%, 50.2%, and 60.1%. Despite the effective reduction in both drag and lift coefficients achieved by the DRL-based AFC strategy, they have not yet stabilized compared to cases with larger $Ar$ values. The initial phase of control exhibits distinct fluctuations, followed by low-frequency and low-amplitude oscillations in the drag and lift coefficients, which are lower than the baseline. Particularly, for $Ar$ values of 0.1 and 1, the drag and lift coefficients experience a sharp decrease initially, followed by low-amplitude, low-frequency periodic oscillations. This indicates that for elliptical cylinders with $Ar$ ranging from 0.5 to 0, the AFC control strategy can partially suppress lift and reduce drag. Correspondingly, for elliptical cylinders with $Ar$ values of 0.5, 0.25, 0.1, and 0, the ratios of the mass flow rates of the synthetic jets relative to the incoming flow rate are 4.9%, 12.6%, 49.9%, and 52.1%, respectively. This suggests that despite the use of higher external energy, the lift coefficient cannot be completely controlled to prevent oscillations and fluctuations. From the vortex plots of the control process, for $Ar$ values of 0.5 and 0.25, the recirculation region behind the elliptical cylinder is elongated during the initial phase of control, with symmetric recirculation bubbles. As the control progresses, the recirculation region becomes unstable, exhibiting lateral oscillations and the shedding of long, narrow vortices. Similarly, for $Ar$ values of 0.5 and 0.25, the recirculation region undergoes elongation during the control phase, with high-frequency and large-amplitude oscillations, accompanied by the shedding of new vortices. After control cessation, the shedding of vortices is somewhat alleviated, with a decrease in both the frequency and intensity of vortex shedding.
In conclusion, for elliptical cylinders with $Ar$ ranging from 1.75 to 0.75, intelligent agents trained using the PPO algorithm have adeptly learned to adopt an energy-efficient flow control strategy. The control policies based on the PPO algorithm significantly reduce lift around the elliptical cylinder, greatly diminish drag, and suppress vortex shedding in the wake. Regarding elliptical cylinder with $Ar$ values ranging from 0.5 to 0, the control strategies obtained from agents trained with the PPO algorithm can moderately alleviate vortex shedding and reduce the amplitude and frequency of fluctuations in lift and drag coefficients. These control outcomes demonstrate several unique capabilities of AFC strategies based on DRL, making them particularly suitable for managing complex fluid dynamics systems. These capabilities include adaptability, real-time optimization, robustness to uncertainties, the ability to handle complex problems, minimal human intervention, and high control efficiency.
Acknowledgements.
The authors would like to express their gratitude to Dr. Jean Rabault (University of Oslo, Oslo, Norway), Dr. Jichao Li (National University of Singapore, Singapore), and Mr. Qiulei Wang (The University of Hong Kong, Hong Kong SAR, China) for making their open-source codes for deep reinforcement learning and numerical simulation available online. These resources have significantly contributed to the research presented in this paper. The relevant code repositories can be accessed at https://github.com/jerabaul29/Cylinder2DFlowControlDRLParallel Rabault and Kuhnle (2019b); Rabault et al. (2019), https://github.com/npuljc/RL_control_Nek5000 Li (2021); Li and Zhang (2022), and https://github.com/venturi123/DRLinFluids Qiulei Wang (2022); Wang et al. (2022b).
Declaration of interests
The authors report no conflict of interest.
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Author ORCIDs